OpenStudy (anonymous):

Solve by factoring: 2a^2+1=-3a

OpenStudy (anonymous):

2a^2+3a + 1 = 0 ( a + 1) ( 2a + 1) = 0 -> a = -1, a = -1/2

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

My pleasure :)